OpenStudy (anonymous):

True or false: The inequality x ≥ -3 is satisfied by point (0, -2).

OpenStudy (anonymous):

since \(0>-3\) the answer is "yes"

OpenStudy (anonymous):

So true?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

Thank you very much:) would you be willing to help me with two more true or false questions?

OpenStudy (anonymous):

sure if they are quick like that one

OpenStudy (anonymous):

True or false: The inequality 2x - 3y ≥ 5 is satisfied by point (1/2, 1/3 ).

OpenStudy (anonymous):

replace \(x\) by \(\frac{1}{2}\) and \(y\) by \(\frac{1}{3}\) and get \[2\times \frac{1}{2}-3\times \frac{1}{3}=1-1=0\] and since \(0\) is NOT greater than \(5\) the answer is FALSE

OpenStudy (anonymous):

True or false: The inequality 5x - 2y + 1 > 0 is satisfied by point (-2, -1).

OpenStudy (anonymous):

repeat as above \[5\times (-2)-2\times (-1)+1\] \[-10+2+1=-8+1=-7\]

OpenStudy (anonymous):

\(-7>0\) is FALSE